OpenStudy (anonymous):

Find the intercepts of the line that contains each pair of points. 1: (-1,-4) and (6,10) 2: (3,4) and (-6,16)

OpenStudy (anonymous):

\[(y-y_{1}) = m(x-x_{1}) \] \[m = \frac{\Delta y}{\Delta x} = \frac{y_{2} - y_{1}}{x_{2} - x_{1}}\] For 1: \[y - 10 = \frac{-4 -10}{-1 - 6}(x - 6)\]\[y - 10 = 2x - 12\]\[y = 2x - 2\] So the y intercept is found by plugging in 0 for x and solving for y: \[y = 2(0) - 2 \rightarrow y = -2\] and the x intercept is found by substituting 0 for y and solving for x: \[0 = 2x - 2 \rightarrow x = 1\]

OpenStudy (anonymous):

Oh, thanks. :)